OpenStudy (anonymous):

Choose the correct simplification of the expression 9/h-3

OpenStudy (abb0t):

It's already in simplest form.

OpenStudy (abb0t):

Unless you mean: \[\frac{ 9 }{ h }-3\] in which case you would find the common denominator.

OpenStudy (anonymous):

go away @abb0t you pissed me off t(-.-t)

OpenStudy (anonymous):

you say choose in the question. are there choices.

OpenStudy (anonymous):

yea.

OpenStudy (abb0t):

LOOOOOOOOOOOOOOOOOOOOOLOLOLOLOLOL

OpenStudy (anonymous):

May I ask what they are?

OpenStudy (anonymous):

But ur cute so i wont hate foever

OpenStudy (anonymous):

h^3/9 9/h^3 9h^3 the expression is already simplified

OpenStudy (anonymous):

is it h-3 or \[h ^{-3}\] in the question

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

then the simplification is \[9h ^{3}\]

OpenStudy (anonymous):

^-^ i have like 3 more....

OpenStudy (anonymous):

Make sure you use ^ in the question or else nobody knows its an exponent

OpenStudy (anonymous):

Ok I'm cool with that. Go ahead

OpenStudy (anonymous):

the second one \[\frac{ 8b ^{3}}{ c ^{3} }\]

OpenStudy (anonymous):

the fourth one \[a ^{2}b\]

OpenStudy (anonymous):

last one. kinda long

OpenStudy (anonymous):

go ahead i got time I'm on here waiting for people to answer my questions lol

OpenStudy (anonymous):

idk what the expanded form is did your teacher tell you how to expand forms and then simplify them?

OpenStudy (anonymous):

is it something like this \[\frac{ x*x*x*x*x*x }{ x*x }\]

OpenStudy (anonymous):

yea its like that

OpenStudy (anonymous):

ok well i guess if you did that you would just cross out 2 from the top and two from the bottom and you'd get \[x ^{4}\] or\[x*x*x*x\] to simplify without using expanded form you would just take the exponent in the numerator and subtract the exponent in the denominator and that's how you get you're answer \[\frac{ x ^{6} }{ x ^{2}}=x ^{6-2}=x ^{4}\]

OpenStudy (anonymous):

as for the second part you can create whatever problem you like and then just show its simplification for example \[\frac{ x ^{20} }{ x ^{12} }=x ^{20-12}=x ^{8}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

hopefully this is what you're looking for this question was a little more in depth but this is what i believe the question means and what the correct answer should be. Good Luck and feel free to fan if you want and ask me any more questions later.

OpenStudy (anonymous):

Alright thanks! <3